OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -8 - 9t. Find the instantaneous velocity at t = 1 by finding the derivative. Tell me if I did it right or not!

OpenStudy (anonymous):

s(t) = -8 - 9t f(x) = -8 - 9x -8-9(x+h) - (-8-9x) / h -8-9x-9h+8+9x / h -9h/h 9

OpenStudy (anonymous):

whooooops meant -9

OpenStudy (anonymous):

good thing I posted this in the history section and not the math one...